Abstract

Coarse Cohomology for Families

We introduce coarse cohomology for families of metric spaces and develop the properties of this cohomology theory. Foliations provide the primary examples of such families and for these there are de Rham, Čech and Alexander-Spanier versions of this theory, which are all isomorphic. We compute the coarse cohomology for a number of important examples.